# Distributionally robust expected shortfall for convex risks

### v1

@@ 5 Conclusion
The transform of a payoff $f$ needed for distributional robustification measured with optimal transport of measures under quadratic cost, is simple and intuitive if $f$ is a maximum of affine functions. Combining this with the dual representation of Expected Shortfall, it is possible to derive an analytical formula for robust Expected Shortfall, subject to an ambiguity tolerance parameter $\theta$. For a three-asset exposure involving call and put options, distributional robust calculations, subject to a choice of $\theta$, gives a quantitative sense of the distributional risk introduced by these options. More importantly, for the quadratic cost function in optimal transport theory, distributional robust calculations for the Expected Shortfall of convex, piecewise-linear, payoffs are only somewhat more complex than the non-robust calculations.
Appendix
Let $X=\mathbb{R}^{n}$ and let $\|x\|$ denote the Euclidean norm of $x\in X$. Recall that we use the quadratic cost function $c(x,y)=\frac{1}{2}\|x-y\|^{2}$ for $x,y\in X$.
Definition 1.
Let $f:X\to\mathbb{R}$ be a convex, piecewise-affine function, $c$ the quadratic cost function, and $\lambda\geq 0$. Then $f^{\lambda c}:X\to\mathbb{R}\cup\{\infty\}$ is defined by the equation
$f^{\lambda c}(x)=\sup_{y\in X}\{f(y)-\lambda c(x,y)\}.$
Observe that $f^{\lambda c}$ is a real-valued function for $\lambda>0$, under the conditions that we impose on $f$ and $c$.
Our result will be based on a fundamental property of the Legendre transform
$g^{*}:x\mapsto~\sup_{y\in X}\{\langle x,y\rangle-~g(y)\}$
of a function $g$. The relationship between the c-transform (i.e. $\lambda=1$) and the Legendre transform is noted in [8]. We modify it for the $\lambda c$-transform, which is a slightly more general case needed in the theorem.
Lemma 1.
For $\lambda>0$, the $\lambda c$-transform $f^{\lambda c}$ of $f$ is related to the Legendre transform $\psi^{*}$ of $\psi(x):=\frac{1}{2}\|x\|^{2}-\frac{1}{\lambda}f(x)$ via the equation
$f^{\lambda c}(x)=-\frac{\lambda}{2}\|x\|^{2}+\lambda\psi^{*}(x).$
Proof.
Using the definition of $f^{\lambda c}$, expanding, and then the definition of $\psi^{*}$, we obtain
$f^{\lambda c}(x)$ $=$ $\sup_{y\in X}\left\{f(y)-\frac{\lambda}{2}\|x-y\|^{2}\right\}$
$=$ $\sup_{y\in X}\left\{f(y)-\frac{\lambda}{2}\|x\|^{2}+\lambda\langle x,y\rangle-\frac{\lambda}{2}\|y\|^{2}\right\}$
$=$ $-\frac{\lambda}{2}\|x\|^{2}+\lambda\sup_{y\in X}\left\{\langle x,y\rangle-\frac{1}{2}\|y\|^{2}+\frac{1}{\lambda}f\right\}$
$=$ $-\frac{\lambda}{2}\|x\|^{2}+\lambda\psi^{*}(x).$
We omit the details of the following standard calculation of the Legendre transform of a quadratic-affine function $x\mapsto\frac{1}{2}\|x\|^{2}+\langle a,x\rangle+b$, where $a,b\in X$, as $x\mapsto\frac{1}{2}\|x-a\|^{2}-b$.
The $\lambda c$-transform, of a maximum of affine functions, turns out to have the same form, with different vertical intercepts.
Proof of Theorem 1
Consider $\psi(x)=\frac{1}{2}\|x\|^{2}-\frac{1}{\lambda}\displaystyle{\max_{1\leq i\leq n}}\left\{\langle m_{i},x\rangle+c_{i}\right\}=\displaystyle{\min_{1\leq i\leq n}}\left\{\frac{1}{2}\|x\|^{2}-\frac{1}{\lambda}\langle m_{i},x\rangle-\frac{c_{i}}{\lambda}\right\}=\displaystyle{\min_{1\leq i\leq n}}\psi_{i}(x)$ where $\psi_{i}(x)=\frac{1}{2}\|x\|^{2}-\frac{1}{\lambda}\langle m_{i},x\rangle-\frac{c_{i}}{\lambda}$. By the above-mentioned Legendre transform of a quadratic-affine function,
$\psi_{i}^{*}(x)=\frac{1}{2}\|x+\frac{m_{i}}{\lambda}\|^{2}+\frac{c_{i}}{\lambda}=\frac{1}{2}\|x\|^{2}+\frac{1}{\lambda}\langle m_{i},x\rangle+\frac{1}{2}\frac{1}{\lambda^{2}}\|m_{i}\|^{2}+\frac{1}{\lambda}c_{i}.$
By a standard property of the Legendre transform of a minimum of functions,
$\psi^{*}(x)=(\displaystyle{\min_{1\leq i\leq n}}\psi_{i}(x))^{*}=\displaystyle{\max_{1\leq i\leq n}}\psi_{i}^{*}(x).$
Combining this with Lemma 1, $f^{\lambda c}(x)=-\frac{\lambda}{2}\|x\|^{2}+\lambda\psi^{*}(x),$ using the above and then Lemma 1 again, we obtain
$f^{\lambda c}(x)=-\frac{\lambda}{2}\|x\|^{2}+\lambda\displaystyle{\max_{1\leq i\leq n}}\psi_{i}^{*}(x)=\displaystyle{\max_{1\leq i\leq n}}\left\{-\frac{\lambda}{2}\|x\|^{2}+\lambda\psi_{i}^{*}(x)\right\}=\displaystyle{\max_{1\leq i\leq n}}\left\{\langle m_{i},x\rangle+c_{i}+\frac{1}{2\lambda}\|m_{i}\|^{2}\right\}.$
$\square$
@@ Acknowledgement
The work is based on research supported in part by the National Research Foundation of South Africa (Grant Number 146018). The author also acknowledges the hospitality of the TU Delft and the University of Vienna, and discussions with Antonis Papapantoleon and Daniel Bartl.

### v2

@@ 4 Conclusion
The transform of a payoff $f$ needed for distributional robustification measured with optimal transport of measures under quadratic cost, is tractable if $f$ is convex, especially if it is a maximum of a finite number of affine function. Combining this with the dual representation of Expected Shortfall, it is possible to derive an analytical formula for robust Expected Shortfall, subject to an ambiguity tolerance parameter $\theta$. For the quadratic cost function in optimal transport theory, distributional robust calculations for the Expected Shortfall of convex payoffs are not much more complex than the non-robust calculations.
@@ Acknowledgement
The work is based on research supported in part by the National Research Foundation of South Africa (Grant Number 146018).